### v2

# Extreme Points and Large Contests*

Abstract.
In this paper, we characterize the extreme points of a class of multidimensional monotone functions. This result is then applied to large contests, where it provides a useful representation of optimal allocation rules under a broad class of distributional preferences of the contest designer. In contests with complete information, the representation significantly simplifies the characterization of the equilibria.
JEL Codes: C65, D72
Keywords: Extreme Points, Monotone Functions, Large Contests
*I am particularly indebted to Ioan Manolescu for his generous time. I also thank Yaron Azrieli, Christian Chambers, Christian Ewerhart, Dominik Inauen, Byeong-hyeon Jeong, Patrick Lahr, Igor Letina, Brendan Pass, Philipp Strack and Stefan Wenger for helpful discussions.
University of Fribourg, Department of Economics. Email: giovanni.valvassoribolge@unifr.ch
1. Introduction
The theory of contests has provided a unifying framework to model a wide range of economic environments. [^1] For a thorough review, see Konrad (2009) Although the classic contributions consider a setting with a finite number of agents [^2] See, inter alia, Lazear and Rosen (1981) and Tullock (2001)., some recent contributions analyze contests with a continuum of agents. [^3] See Olszewski and Siegel (2016), Adda and Ottaviani (2024) and Azrieli and Chambers (2024) among others.
In this paper, we consider the optimal contest design for a broad class of distributional preferences, and we are interested in characterizing the optimal allocation rule, namely the mapping from the (possibly noisy) performance of the agents to the prizes. We show that the best the contest designer can do is to set a two-threshold scheme: the optimal allocation rule is a step function with at most two steps.
This model provides a framework to rationalize several real-world examples of large contests, whose design involve assigning prizes through a combination of a random and deterministic selection procedure. Applications span college and university admissions, as well as funding schemes by private and public institutions.
For example, the Swiss National Science Foundation (SNSF) evaluates funding decisions for grants by selecting a threshold performance (‘funding line’); grant proposals that place above the funding line will obtain funding, whereas those below the funding line will not receive any. Moreover, if multiple proposals place themselves at the funding line, a lottery might be involved to select those who will receive funding. [^4] Details can be found at https://www.snf.ch/en/6cs2wnfJtcfFDL6o/page/evaluation-procedure.
In addition to this, the adoption of lotteries in admission for medical schools is seen as a tool to promote equal opportunity and improve diversity among students. In some European countries, such as the Netherlands and Germany, weighted or unweighted lotteries determine the allocation of some places which remain after a first round of admission in medical school. [^5] For example, the University of Jena recently introduced a lottery (‘losverfahren’) to assign those places that are left out after a first round of selection; see https://jenamedia.de/uni-jena-extra-chance-auf-studienplatz-per-losverfahren. At the University of Amsterdam, the admission cycle for the academic year 2026/2027 in medicine will take the form of an unweighted lottery. According to the Vice Dean of Education at the University of Amsterdam: “Our aim is to promote equal opportunities and diversity among potential students. After all, it gives all candidates an equal chance of being admitted”. See https://student.uva.nl/artikelen/2025-bachelor-geneeskunde-uva-kiest-voor-loting-bij-selectie-van-eerstejaars-2.
For a broad class of distributional preferences of the contest designer, our model provides an economic foundation for two-tier prize schemes, and it encompasses the real-world examples just described.
In order to derive the optimal contest design, we leverage the geometric properties of the set of admissible allocation rules. In particular, we provide a novel characterization of the extreme points of a particular class of multidimensional monotone functions which naturally arises in large contests. Finally, we leverage our result to analyze the large-contest version of a canonical contest with finitely many agents, and we show that the equilibrium analysis significantly simplifies.
The paper is structured as follows. Section 2 introduces the mathematical framework and derives the extreme points result. Section 3 then applies the result to large contests. Section 5 concludes.
Related Literature
This paper contributes to the theory of large contests and to a growing literature on the characterization of extreme points in economics.
Following the pioneering contributions by Olszewski and Siegel (2016) and Olszewski and Siegel (2016), Morgan et al. (2022) and Azrieli and Chambers (2024) have considered contests with a continuum of agents and provided key insights into how this setting significantly simplifies equilibrium analysis with respect to the classic contest models with a finite number of agents. In particular, Azrieli and Chambers (2024) provide an axiomatization of a class of Contest Success Functions (CSFs) in large contests. They show that any CSF satisfying a set of axioms arises as a result of the following steps: the performance of an agent is randomly mapped into a performance; the contest designer sets a performance threshold so as to satisfy a budget constraint; the agents whose performances exceed this cutoff win a prize. They refer to this class of CSFs as Random Performance Functions (RPFs). We contribute to this literature by considering the optimal contest design by a principal who is able to design the allocation rule. We show that RPFs are optimal under specific distributional goals of the contest designer. To the best of my knowledge, optimal contest design has been addressed only by Letina et al. (2023) in a setting with finitely many agents.
A recent literature systematically applies the study of extreme points of monotone functions to several economic settings; e.g., Kleiner et al. (2021), Yang and Zentefis (2024) and Yang and Yang (2025). These papers consider settings in which the revelation principle applies, and hence the class of monotone functions studied maps the (possibly multidimensional) types of the agents into a payoff. In this paper, we consider allocation rules in contests, namely functions which are monotone in agents’ effort, but which also depend on the overall distribution of effort exerted by all the other agents. Hence, the results from the papers cited above are not readily applicable, and a new result on extreme points of this class of monotone function is needed.
A similar model is considered in Lagziel and Lehrer (2022). However, importantly in this paper we consider a large contest environment. Therefore, we can leverage a law-of-large-number type of argument to establish the optimality of threshold strategies, which are not necessarily optimal in Lagziel and Lehrer (2022).
From a methodological perspective, a closely related paper is Jeong and Park (2025). They consider a mechanism design setting and they derive the optimality of a two-tier quota mechanism, which is analogous to the one derived in this paper. However, the economic environment is fundamentally different, and it is not related to contests.
2. Extreme Points
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a compact set and denote $\Delta(\mathcal{X})$ the set of Borel measures on $\mathcal{X}$.
Let $\mathcal{F}$ be the set of monotone functions
$\mathcal{F}=\left\{f:\mathcal{X}\times\Delta(\mathcal{X})\longrightarrow[0,1]\ ,\ f(\cdot,\mu)\ \text{increasing}\ \text{for every}\ \mu\in\Delta(\mathcal{X})\right\}\ .$
We characterize the extreme points of the set of functions
$\mathcal{F}^{*}=\left\{f\in\mathcal{F}\ :\int_{\mathcal{X}}f(x,\mu)\ d\mu(x)\leq k,\ k\in(0,1)\right\}\ .$
A result by Winkler (1988) shows that the set of extreme points of $\mathcal{F}^{*}$ is a convex combination of at most two extreme points of $\mathcal{F}$; hence, we show that they are a mixture of at most two indicator functions defined on up-sets. Moreover, by a characterization of Yang and Yang (2025), these two up-sets are ordered by set inclusion.
Theorem 1.
The set of extreme points of $\mathcal{F}^{*}$ is given by
$\operatorname{ext}(\mathcal{F}^{*})=\left\{f^{*}\in\mathcal{F}^{*}:f^{*}(x,\mu)=\lambda\mathbb{I}_{A_{1}}+(1-\lambda)\mathbb{I}_{A_{2}}\right\}$
for each $\mu\in\Delta(\mathcal{X})$ and some $A_{1}\subseteq A_{2}$.
In other words, extreme points in $\mathcal{F}^{*}$ are step functions on $\mathcal{X}$ with at most two jumps.
Maximizing a Functional
Consider the case $\mathcal{X}=[\underline{x},\overline{x}]$. [^6] For notational convenience, in this subsection we assume that $\mathcal{X}=[0,1]$. We are interested in maximizing a functional $\Pi:\mathcal{F}^{*}\longrightarrow\mathbb{R}$; formally, we are interested in the following problem
$\max_{f\in\mathcal{F}^{*}}\ \Pi(f)=\int_{\mathcal{X}}\Phi\left(f(x,\mu\right),x)\ d\mu(x)\ .$ (1)
As in Kleiner et al. (2021), define the majorization relation $\succ$ as follows.
Definition 1.
For each $\mu\in\Delta(\mathcal{X})$ and for any two functions $f,g\in\mathcal{F}$, we say that $f$ majorizes $g$, $f\succ g$ if
$\int_{t}^{1}f(t,\mu)\ d\mu(t)\geq\int_{t}^{1}g(t,\mu)\ d\mu(t)\ .$
An elegant result due to Fan and Lorentz (1954) gives necessary and sufficient conditions for a functional to be monotonic in the majorization order.
Theorem 2 (Fan and Lorentz (1954)).
Let $\Phi:\mathcal{X}\times\Delta(\mathcal{X})\times\mathcal{X}\longrightarrow[0,1]$. Then,
$\int_{X}\Phi(f(x,\mu),x)\ d\mu(x)\geq\int_{X}\Phi(g(x,\mu),x)\ d\mu(x)$
holds for any two functions $f,g\in\mathcal{F}$ such that $f\succ g$ if and only if $\Phi(u,t)$ is convex in $u$ and supermodular in $(u,t)$.
By Bauer’s Maximum Principle, we know that if $\Phi$ is convex in $f$, the maximizers of (1) can be found among the extreme points of the feasible set. Moreover, the set of extreme points $\mathcal{F}^{*}$ is naturally ordered with respect to the majorization order.
Theorems 1 and 2 yield simple solutions to (1). In the next section, we show the usefulness of these results in the framework of large contests.
3. Large Contests
There is a measure-one mass of agents. They can choose an effort level $x\in\mathcal{X}=[\underline{x},\overline{x}]$. Define $\Delta(\mathcal{X})$ as the set of probability measures on $\mathcal{X}$. A contest designer needs to choose an Allocation Rule; i.e., a mapping $q:\mathcal{X}\times\Delta(\mathcal{X})\longrightarrow[0,1]$. Following Azrieli and Chambers (2024), we assume the designer has a budget sufficient to give prizes to a mass $k\in(0,1)$ of agents. Denote $\mu\in\Delta(\mathcal{X})$ the potential distribution of efforts (the ‘competition’) faced by an agent with effort $x$.
Definition 2.
For every $x\in\mathcal{X}$ and $\mu\in\Delta(\mathcal{X})$, an Allocation Rule $q:\mathcal{X}\times\Delta(\mathcal{X})\longrightarrow[0,1]$ is non-decreasing in $x$ and satisfies
$\int q(x,\mu)d\mu(x)=k.$ (2)
The interpretation of (2) is the following: For any distribution of efforts induced by a strategy profile, the total mass of winners will be equal to $k$. In other words, the prizes will be allocated, regardless of the equilibrium distribution of efforts. [^7] All results go through even in the case where the designer can choose not to allocate a share of the prizes.
3.1. Complete Information
Consider the case of perfectly observable effort. Exerting effort has a cost $c(x)$, where $c(\cdot)$ is a convex function. Winning a prize $V$ yields utility $u(V)$, where $u(\cdot)$ is a continuous and non-decreasing function. Agents’ payoff function is given by
$U(x)=q(x,\mu)u(V)-c(x)\ .$ (3)
In this setting, $q(x,\mu)$ is a CSF [^8] When effort is perfectly observable, the distinction between Allocation Rule and CSF is immaterial.; i.e., for a given competition $\mu\in\Delta(\mathcal{X})$, it specifies the probability of winning a prize for an agent exerting effort $x$.
Denote $\mathbf{M}$ for a probability measure on $(\Delta(\mathcal{X}),\mathcal{B}(\Delta(\mathcal{X})))$. [^9] Endowed with the topology of weak convergence, $\Delta(\mathcal{X})$ has a natural notion of $\sigma$-algebra. Formally, $\mu_{n}\rightarrow\mu\iff\int_{\mathcal{X}}gd\mu_{n}\rightarrow\int_{\mathcal{X}}gd\mu\$, for any $g$ continuous. The payoff of the designer is given by
$\Pi=\int_{\Delta(\mathcal{X})}\int_{\mathcal{X}}q(x^{*},\mu)\pi(x^{*})\ d\mu(x^{*})\ d\mathbf{M}(\mu)\ ,$ (4)
where $x^{*}$ maximizes (3), and $\pi(\cdot)$ is a continuous function.
The specification in (4) is consistent with very different distributional preferences of the designer, depending on the the monotonicity of $\pi(x)$.
Suppose that $\pi(x)$ is increasing; intuitively, the designer cares relatively more about the better-performing agents. As an example, consider a private foundation sponsoring fellowships for students to pursue graduate studies. It might be argued that the foundation is interested in awarding these fellowships to the best-performing students [^10] Of course, if performance is noisy, a best-performing student might not necessarily be one of the students who exerted the most effort. However, by a law-of-large-number type of argument (ensured by the weak convergence of measures), on average it will be the case that better-performing students will be those who have exerted more effort., as it has a stake in their future success, but not necessarily in all the other applicants’.
On the other hand, suppose that $\pi(x)$ is decreasing. This specification might reflect the designer’s concerns about equal opportunity among competing agents. Alternatively, the designer might be concerned that agents have to exert extremely costly (or even wasteful) effort. [^11] See Olszewski and Siegel (2019) and the discussion therein about cram schools in Korea.
The designer chooses an Allocation Rule satisfying (2) in order to maximize (4). The following result readily follows from Theorems 1 and 2.
Theorem 3.
For any $\mathbf{M}$, the set of maximizers is such that
  • If $\pi(x)$ is increasing, the optimal Allocation Rule is given by
[…]
    where, for each $\mu\in\Delta(\mathcal{X})$, $\tilde{x}\in\mathcal{X}$ is the unique solution to (2);
  • If $\pi(x)$ is decreasing, the optimal Allocation Rule is given by $q^{*}(x,\mu)=\{\bar{q}\}$, where $\bar{q}(x,\mu)=k$, for any $x\in\mathcal{X}$ and $\mu\in\Delta(\mathcal{X})$;
  • If $\pi(x)$ is non-monotonic, the optimal Allocation Rule is given by
    $q^{*}(\cdot,\mu)=\begin{cases}&1\qquad x>x^{\prime\prime}\\ &\alpha\qquad x^{\prime}\leq x\leq x^{\prime\prime}\\ &0\qquad x<x^{\prime}\ ,\end{cases}$ (6)
    for some $x^{\prime},x^{\prime\prime}\in X$ and $\alpha\in[0,1]$.
Theorem 3 characterizes the optimal allocation rules. It provides an economic foundation for the allocation rules discussed in the Introduction. Moreover, Theorem 3 significantly simplifies the characterization of equilibria in this class of games with complete information. To this end, consider the case in which $\pi(x)$ is increasing. Then, it is straightforward to derive the following result.
[…]
Figure 1. An optimal allocation rule with two steps.
Proposition 1.
Suppose $\pi(x)$ is increasing. Then, the (essentially) unique equilibrium is for a share $k$ of agent to exert maximal effort $x^{*}=\overline{x}$; and for the remaining agents to stay out of the contest.
Proof.
Define $\bar{x}$ as $u(V)=c(\bar{x})$. Suppose there is an interior equilibrium effort level $\tilde{x}<\overline{x}$. Then, from (2) we know that
$\int_{\mathcal{X}}q^{*}(x,\delta_{\tilde{x}})\ d\delta_{\tilde{x}}(x)$ $=$
[…]
Hence, the optimal $q^{*}$ is given by
[…]
But then, increasing the effort to $\tilde{x}+\epsilon$ yields a discontinuous jump in payoff, and therefore $\tilde{x}$ cannot constitute an equilibrium effort level. Moreover, all agents exerting maximial effort cannot constitute an equilibrium either. In fact,
$U(x)$ $=q^{*}(\bar{x},\delta_{\bar{x}})u(V)-c(\bar{x})$
$=ku(V)-c(\bar{x})$
$<u(V)-c(\bar{x})$
$=0$
Then, the only equilibrium can thus be $\overline{x}$ for a share $k$ of agents, whereas the remaining 1k1-k stay out of the contest. All agents earn zero payoff. ∎
Intuitively, the discontinuity at the threshold value $\tilde{x}$ in (7) gives agents an incentive to slightly increase effort so as to gain a discontinuous jump in the probability of winning. Therefore, the contest designer induces the agents to increase effort until their utility gets to zero: either they stay out, or they exert maximal effort.
Similarly, when $\pi(x)$ is non-monotone, the designer might prefer agents to exert some intermediate level of effort. Then, by setting the two thresholds appropriately, the designer can induce agents to do so by extending the jump in the middle, so as not to give them the incentive to slightly increase their effort level, for any given competition $\mu$ they face.
3.2. Incomplete Information
Suppose now that performance is noisy. In particular, for any $\mu\in\Delta(\mathcal{X})$, each effort level induces a distribution $F(x,\mu)$ of performance. We assume $F(\cdot,\mu)$ to be continuous and non-decreasing. Agents’ payoffs are now given by
$U(x)=F(x,\mu)u(V)-c(x).$ (8)
With a slight abuse of notation, we refer to $F_{q}(x,\mu)$ as a Contest Success Function induced by the Allocation Rule $q$. Clearly, for any two allocation rules $q^{\prime}$ and $q^{\prime\prime}$, it holds that $F_{q^{\prime}}\succ F_{q^{\prime\prime}}$ if and only if $q^{\prime}\succ q^{\prime\prime}$. Therefore, we have the following result.
Corollary 1.
Theorem 3 still holds under incomplete information.
Corollary 1 provides a theoretical foundation for the Random Performance Functions axiomatized in Azrieli and Chambers (2024): If $\pi(x)$ is increasing, the optimal allocation rule induces a CSF which takes the form of a RPF.
4. Extensions
4.1. More General Objective Function
A common specification for the designer’s payoff is given by
$\Pi^{\prime}=\int_{\mathcal{X}}\pi^{\prime}(x)\ d\mu(x)\ ,$ (9)
where $\pi^{\prime}(x)$ is a continuous function.
In other words, the designer might be interested in maximizing some aggregate measure of effort, and s/he might also be concerned with the agents who do not win a prize. The result derived in Theorem 3 goes through also for the specification in (9), both when agents utilities are given by (3) and (8).
Lemma 1.
The Allocation Rule which maximizes (9) subject to (2) is the same as in Theorem 3.
Proof.
Lemma 6 still applies to this specification. Moreover, $q\mapsto\Pi^{\prime}(q)$ is trivially convex in $q$. Bauer’s Maximum Principle then applies as before. ∎
Interestingly, the Allocation Rule which is assumed in Morgan et al. (2022) is not optimal; given that the objective function is non-monotone in effort, a two-tier Allocation Rule yields a higher expected payoff for the designer.
4.2. Multiple Prize Levels
Suppose there are different prize levels $\mathbf{k}=\{k_{1},k_{2},\dots k_{n}\}$, such that $\sum_{i}^{n}k_{i}=k\in(0,1)$, with respective values $V_{1},V_{2},\dots,V_{n}$. The following result shows that it is without loss of generality for the contest designer to focus on a single prize.
Lemma 2.
$k\mapsto\Pi(k)$ is concave.
Proof.
Denote $\mathcal{F}^{o}(k)$ as the set of maximizers. Since $\mathcal{F}^{*}$ is convex, we have that, for $f_{1}\in\mathcal{F}^{o}(k_{1})$ and $f_{2}\in\mathcal{F}^{o}(k_{2})$, and any $k_{1},k_{2}\in(0,1)$
$\Pi(\lambda k_{1}+(1-\lambda)k_{2})$ $\geq\Pi(\lambda f_{1}+(1-\lambda)f_{2})$
$=\lambda\Pi(f_{1})+(1-\lambda)\Pi(f_{2})$
$=\lambda\Pi(k_{1})+(1-\lambda)\Pi(k_{2})\ .$
4.3. Large Contests with Diversity Constraints
Consider now the following variation of the model. In addition to choosing an effort level $x\in\mathcal{X}$, suppose agents have an observable characteristic $\theta\in\Theta=\{\theta_{1},\theta_{2}\}$, with relative shares $\alpha_{1},\alpha_{2}$, respectively. Assume the contest designer faces an additional constraint to give a share $h$ of the prizes to the subset of agents with characteristic $\alpha_{1}$:
$\int_{\mathcal{X}}q(x,\mu_{|\theta_{1}})\ d\mu_{|\theta_{1}}(x)=h\ ,$
with $k>\alpha_{1}>h$.
This additional constraint might be interpreted as a diversity constraint, or a quota on the number of agents with a given characteristic.
For example, at EPFL in Switzerland, the number of students admitted to the Bachelor’s programmes has recently been capped at 3000: Swiss students holding the relevant qualifications will be admitted without limitations, whereas foreign students will be admitted according to the number of remaining places. [^12] See https://actu.epfl.ch/news/admissions-to-bachelor-programs-will-be-limited-st/.
The payoff of the designer can now be written as
$\Pi$ $=\int_{\Delta(\mathcal{X})}\int_{\mathcal{X}}q(x,\mu)\pi(x)\ d\mu(x)\ d\mathbf{M}(\mu)$
$=\int_{\Delta(\mathcal{X})}\alpha_{1}\int_{\mathcal{X}}q(x,\mu_{|\theta_{1}})\pi(x)\ d\mu_{|\theta_{1}}(x)+\alpha_{2}\int_{\mathcal{X}}q(x,\mu_{|\theta_{2}})\pi(x)\ d\mu_{|\theta_{2}}(x)\ d\mathbf{M}(\mu)$
$=\int_{\Delta(\mathcal{X})}\alpha_{1}\Pi(\mu_{|\theta_{1}})+\alpha_{2}\Pi(\mu_{|\theta_{2}})\ d\mathbf{M}(\mu)\ .$
The designer’s problem is then equivalent to two subproblems, one for each subset of agents $\mu|_{\theta_{i}}$, $i=1,2$.
5. Conclusion
This paper characterizes the set of extreme points of a class of multidimensional monotone functions and applies it to large contests. Several directions for future research remain to be pursued. First, from a mathematical perspective the results hinge on the assumption that the effort level has compact support. Relaxing this assumption bears economic interest, although it would significantly complicate the analysis.
Second, the results strongly follow from the additive separable specification of the agents’ utility function in (3) and (8). Assuming a more general functional form might preclude the possibility to apply Bauer’s Maximum Principle. Understanding under which conditions the results are robust to a more general specification is an interesting endeavour, and it is left for future research.
Appendix A Proofs
A.1. Proof of Theorem 1
Proof.
Define $\mathcal{F^{\prime}}=\left\{f:\mathcal{X}\longrightarrow[0,1],f\ \text{increasing}\right\}$.
Step 1.
The set of extreme points of $\mathcal{F}$.
By Choquet (1954), the set of extreme points of $\mathcal{F}^{\prime}$ is given by the functions $f=\mathbb{I}_{[a,1]}$ (or $g=\mathbb{I}_{(a,1]}$), where $\mathbb{I}_{A}$ denotes the indicator function on $A$. The following lemma relates the extreme points of $\mathcal{F}^{\prime}$ and $\mathcal{F}$.
Lemma 3.
Suppose $f$ is extremal in $\mathcal{F}$. Then, $f(\cdot,\mu$) is extremal in $\mathcal{F^{\prime}}$.
Proof.
 [^13] I thank Stefan Wenger for suggesting this beautiful proof.
Let $f$ be extremal in $\mathcal{F}$. Suppose that $\exists\mu_{0}$ such that $f(\cdot,\mu_{0})$ is not extremal in $\mathcal{F^{\prime}}$. Then, there exist $g_{0}$, $g_{1}\in\mathcal{F^{\prime}}$ and $t\in(0,1)$ such that $f(\cdot,\mu_{0})=tg_{0}+(1-t)g_{1}$. Define
$f_{i}(\cdot,\mu)=\begin{cases}&f(\cdot,\mu)\qquad\mu\neq\mu_{0}\\ &g_{i}(\cdot)\qquad\ \ \mu=\mu_{0}\end{cases}$ (10)
for $i=0,1$. But then we have that $f_{0}\neq f_{1}$ and $f=tf_{0}+(1-t)f_{1}$. Hence, $f$ is not extremal. ∎
A theorem in Winkler (1988) shows that an extreme point of $\mathcal{F}^{*}$ is a convex combination of at most two extreme points of $\mathcal{F}$. Hence, extreme points of $\mathcal{F}^{*}$ are of the form
$f^{*}=\lambda\mathbb{I}_{A_{1}}+(1-\lambda)\mathbb{I}_{A_{2}}\ .$
Finally, Proposition 3.1 in Yang and Yang (2025) gives a characterization of monotone functions as probability distributions over ‘nested’ up-sets; i.e., up-sets ordered by set inclusion. Therefore, $A_{1}\subseteq A_{2}$. This concludes the proof of Theorem 1. ∎
A.2. Proof of Theorem 3.
Proof.
Several steps are involved.
Step 2.
$\mathcal{F}$ is a subset of a vector space.
Define
$\mathcal{H}=\left\{f:\mathcal{X}\times\mathcal{M}_{\text{signed}}\longrightarrow\mathbb{R}\ ,\ f\ \text{bounded}\right\}\ ,$
where $\mathcal{M}_{\text{signed}}$ is the set of finite signed measures. We give it the linear structure. Then, since $\Delta(\mathcal{X})$ is a subset of $\mathcal{M}_{\text{signed}}$, it follows that $\mathcal{F}\subseteq\mathcal{H}$. Hence, $\mathcal{F}$ is a subset of a vector space.
Step 3.
$\mathcal{F}$ is compact and convex.
Convexity readily follows because the convex combination of elements in $\mathcal{F}$ is increasing. Compactness requires some functional analytic tools.
Lemma 4.
$\mathcal{F}$ is compact with respect to the product topology.
Proof.
$\mathcal{F}$ can be rewritten as
$\mathcal{F}$ $=\{\psi:\Delta(\mathcal{X})\rightarrow\mathcal{F^{\prime}}\}$
$=\prod_{\mu\in\Delta(\mathcal{X})}\mathcal{F}^{\prime}\ .$
Since each $f\in\mathcal{F}^{\prime}$ is bounded and uniformly integrable, by Dunford-Pettis Theorem $\mathcal{F}^{\prime}$ is compact with respect to the weak topology on $L^{1}(\mu)$, for each $\mu\in\Delta(\mathcal{X})$. Moreover, $\Delta(\mathcal{X})$ is compact with respect to the weak topology.
Tychonoff’s Theorem implies that $\mathcal{F}$ is compact in the product topology of weak topologies. ∎
Step 4.
Bauer’s Maximum Principle.
It is straightforward to show that $\mathcal{F}^{*}$ is also compact and convex. Convexity follows form the linearity of the constraints. Moreover, since $\mathcal{F}^{*}$ is closed and $\mathcal{F}^{*}\subseteq\mathcal{F}$, then $\mathcal{F}^{*}$ is compact.
We then apply the following theorem.
Theorem 0 (Bauer’s Maximum Principle Ball (2023)).
Let $A$ be a locally convex, topological vector space, $B$ be a non-empty compact, convex subset of $A$, and $S:B\longrightarrow\mathbb{R}$ be a quasiconvex function. Then the set $E$ of maximizers of $S$ over $B$ is a face of $B$. Furthermore, $E$ contains an extreme point of $B$.
Clearly, (1) is continuous with respect to the weak topology on $L^{1}(\mu)$ for each $\mu\in\Delta(\mathcal{X})$; hence, it is continuous with respect to the product topology of weak topologies. Therefore, we know that maximizers of (1) are of the form of Theorem 1.
The following results shows that maximizers of (1) are also maximizers of (4).
Lemma 5.
$q^{*}$ is independent of $\mathbf{M}$.
Proof.
Finally, define $q^{*}$ as in (5), for all $\mu\in\Delta(\mathcal{X})$. Then, for any $q\in\mathcal{Q}$, we have
$\Pi(q)=\int_{\mathcal{M}_{\text{prob }}}\Pi(q\mid\mu)d\mathrm{M}(\mu)$ $\leq\int_{\mathcal{M}_{\text{prob }}}\max_{q^{\prime}\in\mathcal{Q}}\Pi(q^{\prime}\mid\mu)d\mathrm{M}(\mu)$
$=\int_{\mathcal{M}_{\text{prob }}}\Pi\left(q^{*}\mid\mu\right)d\mathrm{M}(\mu)=\Pi\left(q^{*}\right)\ .$
Lastly, in order to apply Theorem 2, we need to show that (4) is convex in $q$ and supermodular in $(q,x)$. We have the following result.
Lemma 6.
$x^{*}(q)$ is non-decreasing. In particular, $dx^{*}(q)/dq=0$.
Proof.
Since $u(V)$ is non-decreasing in $x$, $U(x)$ is supermodular in $(q,x)$. By Topkis’ Theorem, $x^{*}(q)$ is non-decreasing. Moreover, since $d^{2}U(x)/dqdx=0$, it follows from the same theorem that $dx^{*}(q)/dq=0$. ∎
Therefore, it follows that (4) is linear, and hence convex in $q$. Moreover, it is supermodular in $(q,x)$ if and only if $\pi$ is increasing in $x$. The submodular case is analogous.
A.3. Alternative Proof of Theorem 3.
Proof.
We provide an alternative proof of Theorem 3, without using Theorem 2.
Proposition 2.
$\Pi(f^{*})\geq\Pi(f)$, for all $f\in\mathcal{F}^{*}$ if and only if $\pi(x)$ is increasing. $\Pi(\bar{f})\geq\Pi(f)$, for all $f\in\mathcal{F}^{*}$ if and only if $\pi(x)$ is decreasing.
Proof.
We first consider the set of measures admitting density functions. Then, we turn to measures with atoms.
Step 5.
Suppose that $\mu\in\Delta(\mathcal{X})$ admits a density $h_{\mu}$.
Denote $f^{*}=\mathbb{I}_{\left\{x\geq\alpha(k)\right\}}$. Then,
$\Pi(f^{*})=\int_{\alpha(k)}^{\infty}\pi(x)h_{\mu}(x)\ dx\ ,$ (11)
where
$\int_{\alpha(k)}^{1}h_{\mu}(x)\ dx=k\ .$ (12)
Differentiating (11) with respect to $k$ yields
$\Pi^{\prime}(k)$ $=-\alpha^{\prime}(k)h_{\mu}(x)\pi(\alpha(k))$
$=\pi(\alpha(k))\ ,$
where the second equality follows from substituting (12) after differentiating with respect to $k$. We know from Lemma 2 that $\Pi(k)$ is concave. If $\pi(\cdot)$ is increasing, then it must be that $\alpha(k)$ is decreasing in $k$. Moreover, for any $f=\beta\mathbb{I}_{\left\{x\geq\alpha^{\prime}(k)\right\}}+(1-\beta)\mathbb{I}_{\left\{x\geq\alpha^{\prime\prime}(k)\right\}}$, it holds that $\Pi(f^{*})\geq\Pi(f)$.
If $\pi(\cdot)$ is decreasing, then clearly $f^{*}$ cannot be optimal. We claim that the optimal $f$ is identically equal to $k$. We have that
$\Pi(f)=\int_{\alpha^{\prime}(k)}^{\bar{x}}\beta\pi(x)h_{\mu}(x)\ dx+\int_{\alpha^{\prime\prime}(k)}^{\bar{x}}(1-\beta)\pi(x)h_{\mu}(x)\ dx$
with the constraint
$\int_{\alpha^{\prime}(k)}^{\bar{x}}\beta h_{\mu}(x)\ dx+\int_{\alpha^{\prime\prime}(k)}^{\bar{x}}(1-\beta)h_{\mu}(x)\ dx=k\ .$
We consider a mean-preserving variation on the constraint by slightly lowering the first threshold, and by slightly increasing the second threshold:
$\int_{\alpha^{\prime}(k)-\epsilon}^{\bar{x}}\beta h_{\mu}(x)\ dx+\int_{\alpha^{\prime\prime}(k)+\epsilon}^{\bar{x}}(1-\beta)h_{\mu}(x)\ dx=k\ .$
By letting $\epsilon\rightarrow 0$, we obtain
$\frac{\partial}{\partial\epsilon}\Pi(f)\biggr\rvert_{\epsilon=0}$ $=\beta h_{\mu}(x)\pi(\alpha^{\prime}(k))-(1-\beta)h_{\mu}(x)\pi(\alpha^{\prime\prime}(k))$
$=\beta h_{\mu}(x)[\pi(\alpha^{\prime}(k))-\pi(\alpha^{\prime\prime}(k)]$
$\geq 0\ ,$
where the second equality follows from substituting the constraint into the objective function, and the last inequality follows because $\pi$ is decreasing.
Step 6.
Suppose that $\mu\in\Delta(\mathcal{X})$ has atoms.
Suppose that $\mu$ is concentrated in two distinct points, $y$, $z\in\mathcal{X}$: $\mu=\gamma\delta_{y}+(1-\gamma)\delta_{z}$, with $\mu(y)=\mu(z)=\frac{1}{2}$, and $\frac{1}{2}<k<1$. Then, we have the following claim.
Claim 1.
A maximizer $f^{*}$ can have at most a single atom if and only if $\pi(x)$ is increasing. The constant function $\bar{f}$ is the (unique) maximizer if $\pi(x)$ is (strictly) decreasing.
Suppose not. Then, there exists $f=\beta\mathbb{I}_{\left\{x>\alpha^{\prime}(k)\right\}}+(1-\beta)\mathbb{I}_{\left\{x>\alpha^{\prime\prime}(k)\right\}}$ such that $\Pi(f)>\Pi(f^{*})$ with $\pi(x)$ increasing. Suppose that $\alpha^{\prime\prime}(k)>z>y>\alpha^{\prime}(k)$. It must hold that
$\int_{\mathcal{X}}f(x,\mu)\ d\mu(x)=\beta\gamma+\beta(1-\gamma)=\beta=k\ .$
Then,
$\Pi(f)$ $=\int_{X}f(x,\mu)\ d\mu(x)$
$=\beta[\gamma\pi(y)+(1-\gamma)\pi(z)]$
$=k[\gamma\pi(y)+(1-\gamma)\pi(z)]\ .$
Now consider $f^{*}=\begin{cases}&1\qquad\quad x>y\\ &\beta\quad\quad\ \ x=y\\ &0\qquad\ \ \ x<y\end{cases}$ . It must hold that
$\int_{\mathcal{X}}f^{*}(x,\mu)\ d\mu(x)$ $=\beta\gamma+(1-\gamma)=k$
$\iff\beta=\frac{k-(1-\gamma)}{\gamma}\ .$
Then,
$\Pi(f^{*})$ $=\int_{X}f^{*}(x,\mu)\ d\mu(x)$
$=\beta\gamma\pi(y)+(1-\gamma)\pi(z)$
$=[k-(1-\gamma)]\pi(y)+(1-\gamma)\pi(z)\ .$
Hence,
$\Pi(f)>\Pi(f^{*})\iff$ $k[\gamma\pi(y)+(1-\gamma)\pi(z)]>[k-(1-\gamma)]\pi(y)+(1-\gamma)\pi(z)$
$\iff$ $(1-k)(1-\gamma)\pi(y)>(1-k)(1-\gamma)\pi(z)$
$\iff$ $\pi(y)>\pi(z)\ ,$
a contradiction, since $\pi(z)\geq\pi(y)$. An analogous argument shows that, when $\pi(x)$ is decreasing, $\bar{f}$ is a maximizer. ∎
Then, each maximizer either belongs to - or it is equal to an element of $\overline{\mathcal{F}^{*}}$, except on a set of measure zero. This concludes the proof. ∎